The accompanying table contains the overall download and upload speeds in mbps for nine carriers in a particular country. CliThe variance for the upload speeds is (Type an integer or decimal rounded to two decimal places as needed.) The standard deviThe coefficient of variation for the upload speeds is %. (Type an integer or decimal rounded to two decimal places as needed.mbps and the difference between the fastest and The average download speed is mbps and the average upload speed is slowest upCarrier Download Speed 22.7 Upload Speed 13.4 Carrier 1 Carrier 2 24.8 14.4 Carrier 3 20.5 8.7 Carrier 4 16.8 12.0 Carrier 5
Answer #1

For Download speed:

Solution: Х x2 22.7 515.29 24.8 615.04 20.5 420.25 16.8 282.24 10.9 118.81 11.2 125.44 10.4 108.16 6.8 46.24 4.3 18.49 { x == value of (**) 9 +1 th observation 2 = value of 5th observation = 11.2 Mode : In the given data, no observation occurs more(2x)2 Ex2 n Population Standard deviation o = n (128.4) 2 2249.96 - 9 9 12249.96 - 1831.84 9 1418.12 = 9 = 746.4578 = 6.816 o

For Upload Speed:

Solution: Х x2 13.4 179.56 14.4 207.36 8.7 75.69 12 144 6.4 40.96 5.6 31.36 6.4 40.96 3.2 10.24 4.2 17.64 { x = 74.3 x2 = 747= value of () 9 +1th observation 2 = value of 5th observation = = 6.4 Mode : In the given data, the observation 6.4 occurs ma(2x)2 { x2 n Population Standard deviation o = n (74.3) 2 1747.77 - 9 II 9 1747.77 - 613.3878 = 9 1134.3822 = 9 II 14.9314 =

The download speed is left skewed

The upload speed is symmetrical since median= mode

I have given my best to solve your problem. Please like the answer if you are satisfied with it. ?